The average cholesterol content is 180 m deviation of 12 milligrams assume that the variable is normally distributed give a sample of 16 pounds is selected what is the probability that the mean of the sample with the greater than 172 milligrams

1
Expert's answer
2022-06-13T14:26:42-0400
"P(X>172)=1-P(Z\le\dfrac{172-180}{12/\sqrt{16}})"

"=1-P(Z\le-2.6667)\approx0.99617"